OpenStudy (anonymous):

solve for x: x/5 - 18 = -3 a. 3 b. 105 c. -105 d. 75

OpenStudy (texaschic101):

multiply the equation by 5 and then solve for x

OpenStudy (austinl):

First I would add 18 to both sides, then multiply both sides by 5. That should give you x.

OpenStudy (texaschic101):

that works too

OpenStudy (anonymous):

im looking for the answer so i can solve the math on my own :p

OpenStudy (austinl):

Riiiiight..... simple problem. Simple steps. Just do what we said and you will get the answer.

OpenStudy (jack1):

\[\frac x5 - 18 = -3\] \[\frac x5 \times 5 - 18 \times 5 = -3 \times 5\] \[ x - 90 = -15\] can u solve from here?

OpenStudy (ineedserioushelp):

x/5 - 18 = -3 add 18 to both sides x/5 = 15 opposite operation to get x 5 * 15 = ? x = the product of ^

OpenStudy (ineedserioushelp):

If you just want answer it is 75